OpenStudy (anonymous):

differentiate y=x/square root (x^2+1)

OpenStudy (anonymous):

Use the quotient rule. Do you know what that is?

OpenStudy (anonymous):

yes i just wanted to know the answer because i keep getting something different to the answer given

OpenStudy (anonymous):

so i'm not sure where i went wrong

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

1/(x^2+1)

OpenStudy (anonymous):

I think you forgot -x

OpenStudy (anonymous):

and the correct answer according to the solution is 1/(x^2+1)^(3/2)

OpenStudy (anonymous):

I got \[(1/(x^2+1) )-x\]

OpenStudy (anonymous):

The derivative is: \[-\frac{x^2}{\left(1+x^2\right)^{3/2}}+\frac{1}{\sqrt{1+x^2}}=\frac{1}{\left(1+x^2\right)^{3/2}} \]

OpenStudy (anonymous):

robtobey could you show me your working out before reaching what you had come up with in the above post thanks!

OpenStudy (anonymous):

Used Mathematica 8. The request for the derivative from that program was:\[D\left[\frac{x}{\sqrt{1+x^2}},x\right] \] Browse over to WolframAlpha.com and enter: derivative of x/square root (x^2+1) Seclect show steps See the attachment

OpenStudy (anonymous):

Product rule: \[ \dfrac{\mathbb{d}}{\mathbb{d}x}(u\cdot v)=u\cdot \dfrac{\mathbb{d}v}{\mathbb{d}x}+v\cdot \dfrac{\mathbb{d}u}{\mathbb{d}x} \] where u, v are functions of x. The quotient rule is just extra baggage.

OpenStudy (anonymous):

thank you robtobey,romero and INewton!! ^.^

OpenStudy (anonymous):

¬_¬ It was all robtobey.

OpenStudy (anonymous):

but INewton made a lovely conclusion to the question!

OpenStudy (anonymous):

=]